A school sold 200 tickets that totaled $980.00. Students ticket $4.00 and Adult ticket $5.50 . How many adult sold and how many students tickets were sold.
Answer by checkley71(8403)

A+S=200 OR A=200-S
5.5A+4S=980
5.5(200-S)+4S=980
1100-5.5S+4S=980
-1.5S=980-1100
-1.5S=-120
S=-120/-1.5
S=80 IS THE NUMBER OF STUDENT TICKETS SOLD.
200-80=120 IS THE NUMBER OF ADULT TICKETS SOLD.
PROOF
5.5*120+4*80=980
660+320=980
980=980